If Ben invests $5000 at 5% interest per year, how much additional money must he
invest at 6.5% annual interest to ensure that the interest he receives each year is 5.5% of the total amount invested?
$_______

Found 2 solutions by checkley71, ankor@dixie-net.com:
Answer by checkley71(8403)

5000*.05+.065x=.055(5000+x)
250+.065x=275+.055x
.065x-.055x=275-250
.01x=25
x=25/.01
x=2500 amount of additional money to be invested @ 6.5%.
proof
5000*.05+.065*2500=.055(5000+2500)
250+162.5=.055*7500
412.5=412.5

Answer by ankor@dixie-net.com(22740)

If Ben invests $5000 at 5% interest per year, how much additional money must he invest at 6.5% annual interest to ensure that the interest he receives each year is 5.5% of the total amount invested?
:
Let x = amt of additional money he has to invest to accomplish this
:
Here is the equation that we want:
:
Money invested at 5% + Money invested at 6.5% = Total money at 5.5%
:
.05(5000) + .065x = .055(x+5000)
250 + .065x = .055x + 275
.065x - .055x = 275 - 250
.01x = 25
x = 25/.01
x = $2500, additional money to be invested
:
:
Check solution using our original equation:
250 + .065(2500) =
250 + 162.5 = 412.50
:
.055(7500) = 412.50 confirms our solution
:
That was pretty easy, right?